OpenStudy (anonymous):

the demand curve for a product is given by p= 120 - 10q. Find the elasticity of demand E when p= 70. If this price rises by 3 percent calculate the corresponding percentage change in demand

OpenStudy (anonymous):

Want dq/q, the elasticity, for given dp/p q = 12-p/10 dq=dp/10 when p = 70, q = 12-70/10 = 12-7 = 5 dp = .03(70) = 2.1 dq = 0.21 dq/q = 0.21/5 = .042 or 4.2%

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are very welcome.

OpenStudy (anonymous):

how would I solve for elasticity? because the question also wants to know what E is equal to